# Federated LLMs Fine-tuned with Adaptive Importance-Aware LoRA

### v1

@@ III Importance-based Parameters Truncation and Freezing Schemes
In this section, to handle the heterogeneity of client resources, we first introduce an importance-based parameter truncation scheme at the local LLM. To address the issue of performance degradation resulting from the truncation scheme, we subsequently introduce an importance-based parameter freezing scheme at the local LLM.
III-A Decomposed LoRA Rank-1 Matrices Importance Evaluation
@@
In FL environments, clients often have different computational capacities, making it challenging to train all clients’ LoRA parameters at a uniform rank across all clients. To address this, the LoRA parameter product $\mathbf{B}\mathbf{A}$ will be decomposed into smaller, rank-1 matrices. As shown in Fig. 1, we take the LoRA $r_{g}=3$ as an example. The cloud server calculates an importance score for each rank-1 matrix and broadcasts this information to the clients. Clients can then prioritize these rank-1 matrices, training the most important ones first.
III-A1 Matrix Decomposition
@@
The cloud server will decompose the product $\mathbf{B}_{g}\mathbf{A}_{g}$ into rank-1 matrices as
$\mathbf{B}_{g}\mathbf{A}_{g}=\sum_{i=1}^{r}\mathbf{b}_{i}\mathbf{a}_{i},$ (3)
@@
III-A2 Element-wise Importance Calculation
To assess the significance of each element within the vectors $\mathbf{b}_{i}$ and $\mathbf{a}_{i}$, we employ a sensitivity-based metric[10]. This involves computing the gradient of the loss function $L$ with respect to each element and using the product of the element’s value and its gradient as a measure of importance. However, in the context of FL, the global model does not have a direct loss function $L$ to calculate the gradient. Instead, we approximate the gradient by scaling the change in parameter values between consecutive communication rounds using the learning rate. Then, we define a unified importance function $I(w_{ij})$ for any trainable parameter $w_{ij}$ as follows:
@@ III-A2 Element-wise Importance Calculation
where $w_{ij}$ represents each element $b_{ji}$ in $\mathbf{b}_{i}$, with $j\in\{1,\ldots,d\}$, or each element $a_{iq}$ in $\mathbf{a}_{i}$, with $q\in\{1,\ldots,l\}$. The term $\Delta w_{ij}=w_{ij}^{(t)}-w_{ij}^{(t-1)}$ represents the change in the parameter value between two consecutive global updates, and $\eta$ is the learning rate.
@@
First, the smoothed sensitivity $\bar{I}^{(t)}(w_{ij})$ is calculated using an exponential moving average:
@@
where $\beta_{1}$ is a smoothing factor between 0 and 1.
Next, the uncertainty $\bar{U}^{(t)}(w_{ij})$ is quantified to capture the variability in the sensitivity scores:
@@
where $\beta_{2}$ is another smoothing factor between 0 and 1.
@@
By incorporating these smoothing techniques, we enhance the reliability of the importance scores, making them more robust to the stochasticity in FL environments.
III-A3 Aggregation of Importance Scores
To obtain the overall importance score for the rank-1 matrix $\mathbf{b}_{i}\mathbf{a}_{i}$, we sum the smoothed and uncertainty adjusted importance scores for all elements in $\mathbf{b}_{i}$ and $\mathbf{a}_{i}$. This approach ensures that the aggregated score reflects the reliability and significance of each parameter. The overall importance score $S_{i}$ of rank-1 matrix is calculated using:
@@ III-A3 Aggregation of Importance Scores
III-A4 Broadcast Importance Scores
After calculating the importance scores $S$ for each rank-1 matrix, we obtain a list of scores $\left[S_{1},S_{2},\ldots,S_{r}\right]$. This list represents the importance of each rank-1 matrix in the decomposition of $\mathbf{B}\mathbf{A}$. The list of importance scores is then broadcast to all clients.
III-B Importance-based Parameters Truncation Scheme
To accommodate diverse computational capabilities among clients, we allow each client to have a different LoRA rank $r_{k}\in[r_{\text{min}},r_{\text{max}}]$. However, this flexibility introduces the challenge of aligning the cloud server’s global model with the different LoRA ranks of each client’s local model. To address this issue, we propose an importance-based parameter truncation scheme.
III-B1 Truncation Process
@@
Each client $k\in\mathcal{K}$ receives the list of importance scores $[S_{1},S_{2},\ldots,S_{r}]$ from the server. Based on these scores, client $k$ selects the top $r_{k}$ rank-1 matrices as
$\mathcal{I}_{k}=\text{topk}\left([S_{1},S_{2},\ldots,S_{r}],r_{k}\right),$ (9)
where $\mathcal{I}_{k}$ represents the set of indices corresponding to the rank-1 matrices selected by client $k$.
@@
Client $k$ then truncates the low-rank matrices $\mathbf{B}_{g}$ and $\mathbf{A}_{g}$ to align with the local model’s LoRA rank requirements. Specifically, the client retains $r_{k}$ column vectors $\mathbf{b}_{i}$ and $r_{k}$ row vectors $\mathbf{a}_{i}$ corresponding to the highest important scores:
$\mathbf{B}_{k}=\mathbf{B}_{g}[:,\mathcal{I}_{k}],\quad\mathbf{A}_{k}=\mathbf{B}_{g}[\mathcal{I}_{k},:].$ (10)
III-B2 Local Model Update
The truncated low-rank matrices $\mathbf{B}_{k}$ and $\mathbf{A}_{k}$ will be copied to the LoRA parameters of client $k$’s local model. Subsequently, client $k$ trains the local model using its local dataset $\mathcal{D}_{k}$. In the parameters truncation scheme, the local loss function for each client $k$ is defined as
$L_{k}(\mathbf{B}_{k},\mathbf{A}_{k})$ $=\frac{1}{|\mathcal{D}_{k}|}\sum_{\xi\in\mathcal{D}_{k}}\ell\left(\left(\mathbf{B}_{k},\mathbf{A}_{k}\right),\xi\mid\mathbf{W}_{\text{pre}}\right)$
$+\frac{\lambda}{2}\left(\|\mathbf{B}_{k}\|^{2}+\|\mathbf{A}_{k}\|^{2}\right),$ (11)
@@ III-B2 Local Model Update
III-C Importance-based Parameters Freezing Scheme
The truncation process introduced above will inevitably result in performance degradation when distributing the model to clients. To address this issue, we propose a novel importance-based parameters freezing scheme, where both local models and the global model utilize the same maximum LoRA rank, denoted as $r_{k}=r_{g}=r_{\text{max}}$. To accommodate client resource constraints, clients will freeze a portion of the LoRA parameters based on their importance. We denote the freezing ratio as $\alpha_{k}$, which represents the proportion of LoRA parameters that are frozen relative to the total parameters.
III-C1 Freezing Process
In the HAFL framework, each client updates its local LoRA parameters by selectively training a subset of the parameters based on the importance scores received from the cloud server. All LoRA parameters $\Theta_{\text{LoRA}}=\{\mathbf{B}_{g},\mathbf{A}_{g}\}$ broadcasted by the server are replicated to the client’s local model, but only the most important rank-1 matrices are unfrozen and optimized by the client’s local dataset. Based on the freezing ratio $\alpha_{k}$ of each client $k$, $(1-\alpha_{k})r_{\rm max}$ rank-1 matrices will be fine-tuned.
@@ III-C1 Freezing Process
Each client $k\in\mathcal{K}$ receives the list of importance scores $[S_{1},S_{2},\ldots,S_{r}]$ from the server. Based on these scores, client $k$ selects the top $(1-\alpha_{k})r_{\rm max}$ rank-1 matrices to actively train, where $\alpha_{k}$ can vary between clients depending on their computational capacity and resource availability. The selection is performed by choosing the indices corresponding to the highest importance scores:
$\mathcal{I}_{k}=\text{topk}\left([S_{1},S_{2},\ldots,S_{r}],(1-\alpha_{k})r_{\rm max}\right),$ (12)
@@
where $\mathcal{I}_{k}$ is the set of indices of the selected rank-1 matrices for client $k$. The remaining rank-1 matrices’ indices are denoted as $\mathcal{U}_{k}=\{1,2,\ldots,r_{\rm max}\}\setminus\mathcal{I}_{k}$.
III-C2 Local Model Update
@@
Each client $k$ updates its local model by minimizing a loss function $L_{k}$ over its dataset $\mathcal{D}_{k}$. The loss function is defined as
$\quad L_{k}(\mathbf{B}_{k},\mathbf{A}_{k})$
$=\frac{1}{|\mathcal{D}_{k}|}\sum_{\xi\in\mathcal{D}_{k}}\ell\left(\left(\mathbf{B}_{k}^{\mathcal{I}_{k}},\mathbf{A}_{k}^{\mathcal{I}_{k}}\right),\xi\;\middle|\;\mathbf{W}_{\text{pre}},\mathbf{B}_{k}^{\mathcal{U}_{k}},\mathbf{A}_{k}^{\mathcal{U}_{k}}\right)$
$+\frac{\lambda}{2}\left(\|\mathbf{B}_{k}^{\mathcal{I}_{k}}\|^{2}+\|\mathbf{A}_{k}^{\mathcal{I}_{k}}\|^{2}\right),$ (13)
where $|\mathcal{D}_{k}|$ represents the size of this dataset, the loss function $\ell$ evaluates the model’s performance on a data sample $\xi$. The matices $\mathbf{B}_{k}^{\mathcal{I}_{k}}=\mathbf{B}_{k}[:,\mathcal{I}_{k}]$ and $\mathbf{A}_{k}^{\mathcal{I}_{k}}=\mathbf{A}_{k}[\mathcal{I}_{k},:]$ represent the trained LoRA parameter rank-1 matrices. Additionally, $\mathbf{B}_{k}^{\mathcal{U}_{k}}=\mathbf{B}_{k}[:,\mathcal{U}_{k}]$ and $\mathbf{A}_{k}^{\mathcal{U}_{k}}=\mathbf{A}_{k}[\mathcal{U}_{k},:]$ are the untrained frozen LoRA parameter rank-1 matrices.
@@
In the importance-based parameters freezing scheme, clients receive LoRA parameters $\{\mathbf{B}_{g},\mathbf{A}_{g}\}$ from the cloud server and replicate all of them to their local LoRA rank-1 matrices. However, only the rank-1 matrices indexed by $\mathcal{I}_{k}$ are trained and optimized, while those indexed by $\mathcal{U}_{k}$ remain frozen during local training.
III-D Adaptive Global Aggregation at the Cloud

### v2

@@ IV Importance-based Parameter Truncation and Freezing Schemes
In this section, to handle the heterogeneity of client resources, we first introduce an importance-based parameter truncation scheme at the local LLMs. To address the issue of performance degradation resulting from truncation, we subsequently introduce an importance-based parameter freezing scheme at the local LLMs.
IV-A Decomposed LoRA Rank-1 Matrix Importance Evaluation
@@
In practical FL environments, clients often have different computational capacities, making it challenging to train all clients’ LoRA parameters with a uniform rank across all clients. To address this, the LoRA parameter product $\mathbf{B}\mathbf{A}$ is decomposed into smaller, rank-1 matrices. As shown in Fig. 1, we take LoRA with $r_{g}=3$ as an example. The cloud server calculates an importance score for each rank-1 matrix and broadcasts this information to the clients. Clients can then prioritize these rank-1 matrices, training the most important ones first.
IV-A1 Matrix Decomposition
@@
The cloud server decomposes the product $\mathbf{B}_{g}\mathbf{A}_{g}$ into rank-1 matrices as
$\mathbf{B}_{g}\mathbf{A}_{g}=\sum_{i=1}^{r_{g}}\mathbf{b}_{i}\mathbf{a}_{i},$ (20)
@@
IV-A2 Element-wise Importance Calculation
To assess the significance of each element within vectors $\mathbf{b}_{i}$ and $\mathbf{a}_{i}$, we employ a sensitivity-based metric[27]. This involves computing the gradient of the global loss function $L$ with respect to each element and using the product of the element’s value and its gradient as a measure of importance. However, in the context of FL, the global model does not have a direct loss function $L$ to calculate the gradient, as the training data is distributed across multiple clients, and the global model is updated based on aggregated client updates. Instead, we approximate the gradient by scaling the change in parameter values between consecutive communication rounds using the learning rate. Then, we define a unified importance function $I(w_{ij})$ for any trainable parameter $w_{ij}$ as
@@ IV-A2 Element-wise Importance Calculation
where $w_{ij}$ represents element $b_{ji}$ in $\mathbf{b}_{i}$, with $j\in\{1,\ldots,d\}$, or element $a_{iq}$ in $\mathbf{a}_{i}$, with $q\in\{1,\ldots,l\}$. The term $\Delta w_{ij}=w_{ij}^{(t)}-w_{ij}^{(t-1)}$ represents the change in the parameter value between two consecutive global updates, and $\eta$ is the learning rate.
@@
First, the smoothed sensitivity $\bar{I}^{(t)}(w_{ij})$ is calculated using an exponential moving average
@@
where $\beta_{1}\in[0,1]$ is a smoothing factor.
Next, the uncertainty $\bar{U}^{(t)}(w_{ij})$ is quantified to capture the variability in the sensitivity scores
@@
where $\beta_{2}\in[0,1]$ is another smoothing factor.
@@
By incorporating these smoothing techniques, we enhance the reliability of the importance scores, making them more robust to the stochasticity of FL environments.
IV-A3 Aggregation of Importance Scores
To obtain the overall importance score for rank-1 matrix $\mathbf{b}_{i}\mathbf{a}_{i}$, we sum the smoothed and uncertainty adjusted importance scores for all elements of $\mathbf{b}_{i}$ and $\mathbf{a}_{i}$. This approach ensures that the aggregated score reflects the reliability and significance of each parameter. The overall importance score $S_{i}$ of the rank-1 matrix is calculated as
@@ IV-A3 Aggregation of Importance Scores
IV-A4 Broadcast Importance Scores
After calculating the importance scores $S$ for each rank-1 matrix, we obtain a list of scores $\left[S_{1},S_{2},\ldots,S_{r_{g}}\right]$. This list represents the importance of each rank-1 matrix in the decomposition of $\mathbf{B}\mathbf{A}$. The list of importance scores is then broadcast to all clients.
IV-B Importance-based Parameter Truncation Scheme
To accommodate diverse computational capabilities of clients, we allow each client to have a different LoRA rank $r_{k}\in[r_{\text{min}},r_{\text{max}}]$. However, this flexibility introduces the challenge of aligning the cloud server’s global model with the different LoRA ranks of each client’s local model. To address this issue, we propose an importance-based parameter truncation scheme.
IV-B1 Truncation Process
@@
Each client $k\in\mathcal{K}$ receives the list of importance scores $[S_{1},S_{2},\ldots,S_{r_{g}}]$ from the server. Based on these scores, client $k$ selects the top $r_{k}$ rank-1 matrices as
$\mathcal{I}_{k}^{\text{T}}=\text{topk}\left([S_{1},S_{2},\ldots,S_{r_{g}}],r_{k}\right),$ (26)
where $\mathcal{I}_{k}^{\text{T}}$ represents the set of indices corresponding to the rank-1 matrices selected by client $k$.
@@
Client $k$ then truncates the low-rank matrices $\mathbf{B}_{g}$ and $\mathbf{A}_{g}$ to align with the local model’s LoRA rank requirement $r_{k}$. Specifically, the client retains $r_{k}$ column vectors $\mathbf{b}_{i}$ and $r_{k}$ row vectors $\mathbf{a}_{i}$ corresponding to the highest importance scores
$\mathbf{B}_{k}^{\text{T}}=\mathbf{B}_{g}[:,\mathcal{I}_{k}^{\text{T}}],\quad\mathbf{A}_{k}^{\text{T}}=\mathbf{A}_{g}[\mathcal{I}_{k}^{\text{T}},:].$ (27)
IV-B2 Local Model Update
The truncated low-rank matrices $\mathbf{B}_{k}^{\text{T}}$ and $\mathbf{A}_{k}^{\text{T}}$ are copied to the LoRA parameters of client $k$’s local model. Subsequently, client $k$ trains the local model using its local dataset $\mathcal{D}_{k}$. In the parameters truncation scheme, the local loss function for each client $k$ is defined as
$L_{k}^{\text{T}}(\mathbf{B}_{k}^{\text{T}},\mathbf{A}_{k}^{\text{T}})$ $=\frac{1}{|\mathcal{D}_{k}|}\sum_{\xi\in\mathcal{D}_{k}}\ell\left(\left(\mathbf{B}_{k}^{\text{T}},\mathbf{A}_{k}^{\text{T}}\right),\xi\mid\mathbf{W}_{\text{pre}}\right)$
$+\frac{\lambda}{2}\left(\|\mathbf{B}_{k}^{\text{T}}\|^{2}+\|\mathbf{A}_{k}^{\text{T}}\|^{2}\right),$ (28)
@@ IV-B2 Local Model Update
IV-C Importance-based Parameter Freezing Scheme
While the truncation process described above offers the advantage of simple implementation, it inevitably results in performance degradation when distributing the model to clients. To address this issue, we propose a novel importance-based parameter freezing scheme, where both local models and the global model utilize the same maximum LoRA rank, denoted as $r_{k}=r_{g}=r_{\text{max}}$. To accommodate client resource constraints, clients will freeze a portion of the LoRA parameters based on their importance. We denote the freezing ratio as $\alpha_{k}$, which represents the proportion of LoRA parameters that are frozen relative to all parameters.
IV-C1 Freezing Process
In the HAFLQ framework, each client updates its local LoRA parameters by selectively training a subset of the parameters based on the importance scores received from the cloud server. All LoRA parameters $\Theta_{\text{LoRA}}=\{\mathbf{B}_{g},\mathbf{A}_{g}\}$ broadcasted by the server are replicated to the client’s local model, where the local LoRA parameters are initialized as $\mathbf{B}_{k}^{\text{F}}=\mathbf{B}_{g}$ and $\mathbf{A}_{k}^{\text{F}}=\mathbf{A}_{g}$, but only the most important rank-1 matrices are unfrozen and optimized by the client’s local dataset. Based on the freezing ratio $\alpha_{k}$ of each client $k$, $(1-\alpha_{k})r_{\rm max}$ rank-1 matrices will be fine-tuned.
@@ IV-C1 Freezing Process
Each client $k\in\mathcal{K}$ receives the list of importance scores $[S_{1},S_{2},\ldots,S_{r_{g}}]$ from the server. Based on these scores, client $k$ selects the top $(1-\alpha_{k})r_{\rm max}$ rank-1 matrices to actively train, where $\alpha_{k}$ can vary between clients depending on their computational capacity and resource availability. The selection is performed by choosing the indices corresponding to the highest importance scores:
$\mathcal{I}_{k}^{\text{F}}=\text{topk}\left([S_{1},S_{2},\ldots,S_{r_{g}}],(1-\alpha_{k})r_{\rm max}\right),$ (29)
@@
where $\mathcal{I}_{k}^{\text{F}}$ is the set of indices of the selected rank-1 matrices for client $k$. The remaining rank-1 matrices’ indices are denoted as $\mathcal{U}_{k}^{\text{F}}=\{1,2,\ldots,r_{\rm max}\}\setminus\mathcal{I}_{k}^{\text{F}}$.
Freezing Low-Rank Matrices
Client $k$ then freezes the low-rank matrices $\mathbf{B}_{k}^{\text{F}}$ and $\mathbf{A}_{k}^{\text{F}}$ to align with the local model’s LoRA freezing ratio requirements. Specifically, the trained LoRA parameter rank-1 matrices are retained as
$\mathbf{B}_{k}^{\mathcal{I}}=\mathbf{B}_{k}^{\text{F}}[:,\mathcal{I}_{k}^{\text{F}}],\quad\mathbf{A}_{k}^{\mathcal{I}}=\mathbf{A}_{k}^{\text{F}}[\mathcal{I}_{k}^{\text{F}},:].$ (30)
Additionally, the frozen LoRA parameter rank-1 matrices are given by
$\mathbf{B}_{k}^{\mathcal{U}}=\mathbf{B}_{k}^{\text{F}}[:,\mathcal{U}_{k}^{\text{F}}],\quad\mathbf{A}_{k}^{\mathcal{U}}=\mathbf{A}_{k}^{\text{F}}[\mathcal{U}_{k}^{\text{F}},:].$ (31)
IV-C2 Local Model Update
@@
Each client $k$ updates its local model by minimizing a loss function $L_{k}^{\text{F}}$ over its dataset $\mathcal{D}_{k}$. The loss function is defined as
$\quad L_{k}^{\text{F}}(\mathbf{B}_{k}^{\mathcal{I}},\mathbf{A}_{k}^{\mathcal{I}})$
$=\frac{1}{|\mathcal{D}_{k}|}\sum_{\xi\in\mathcal{D}_{k}}\ell\left(\left(\mathbf{B}_{k}^{\mathcal{I}},\mathbf{A}_{k}^{\mathcal{I}}\right),\xi\;\middle|\;\mathbf{W}_{\text{pre}},\mathbf{B}_{k}^{\mathcal{U}},\mathbf{A}_{k}^{\mathcal{U}}\right)$
$+\frac{\lambda}{2}\left(\|\mathbf{B}_{k}^{\mathcal{I}}\|^{2}+\|\mathbf{A}_{k}^{\mathcal{I}}\|^{2}\right).$ (32)
@@
In the importance-based parameters freezing scheme, clients receive LoRA parameters $\{\mathbf{B}_{g},\mathbf{A}_{g}\}$ from the cloud server and replicate all of them to their local LoRA rank-1 matrices. However, only the rank-1 matrices indexed by $\mathcal{I}_{k}^{\text{F}}$ are trained and optimized, while those indexed by $\mathcal{U}_{k}^{\text{F}}$ remain frozen during local training.
V Importance-aware Bandwidth-Adaptive Communication Quantization
In wireless FL systems, communication overhead is constrained by limited bandwidth and channel fading, which directly impact the maximum number of bits that can be transmitted. If all parameters are transmitted in full precision, the bandwidth demand would exceed practical limits. On the other hand, static quantization with uniform precision ignores parameter importance, potentially harming model performance.
To address this challenge, we propose an Importance-Aware Bandwidth-Adaptive Communication Quantization scheme for efficient uploading of LoRA parameters. In this scheme, more important LoRA parameters are prioritized for transmission and are assigned higher precision (i.e., more bits), while less important parameters are transmitted with lower precision. The precision of each parameter is dynamically adjusted to align with the bandwidth budget $M_{k}$ of each client $k$, defined in (12) in Section II, ensuring efficient utilization of the available communication resources. If the bandwidth budget is insufficient to transmit all parameters, the least important parameters may be omitted entirely from transmission.
V-A Importance-Based Priority Transmission
During each communication round, clients upload their locally trained LoRA parameters to the central cloud server. The specific parameters uploaded depend on the scheme employed. In the parameter truncation scheme, each client $k$ uploads the LoRA parameters $\mathbf{B}_{k}^{\text{T}}$ and $\mathbf{A}_{k}^{\text{T}}$, as defined in (27). In the parameter freezing scheme, each client $k$ uploads subsets of these parameters, denoted as $\mathbf{B}_{k}^{\mathcal{I}}$ and $\mathbf{A}_{k}^{\mathcal{I}}$, as defined in (30).
Each LoRA parameter matrix can be decomposed into rank-1 matrices, represented as $\mathbf{B}_{k}^{\text{T}}[:,i]\mathbf{A}_{k}^{\text{T}}[i,:]$ for the truncation scheme, where $i\in[0,r_{k})$, or $\mathbf{B}_{k}^{\mathcal{I}}[:,i]\mathbf{A}_{k}^{\mathcal{I}}[i,:]$ for the freezing scheme, where $i\in[0,(1-\alpha_{k})r_{\text{max}})$. Due to the selection process $\mathcal{I}_{k}^{\text{T}}$ in (26) and $\mathcal{I}_{k}^{\text{F}}$ in (29), the rank-1 matrices for these uploaded parameters are already ordered by decreasing importance. The first rank-1 matrix (corresponding to $i=0$) is the most important one, with subsequent matrices decreasing in importance.
To maximize system performance under limited bandwidth, clients transmit the rank-1 matrices in the order of their importance. Specifically, transmission begins with the most important rank-1 matrix, which corresponds to the first column of $\mathbf{B}_{k}^{\text{T}}$ and the first row of $\mathbf{A}_{k}^{\text{T}}$ in the truncation scheme, or the first column of $\mathbf{B}_{k}^{\mathcal{I}}$ and the first row of $\mathbf{A}_{k}^{\mathcal{I}}$ in the freezing scheme. Subsequent rank-1 matrices are transmitted in sequence, with their importance progressively decreasing. By prioritizing the transmission of the most important parameters, the system ensures that the limited bandwidth is utilized to maximize the overall performance of the FL system.
V-B Importance-Aware Bandwidth-Adaptive Communication Quantization
Building on the importance-based priority transmission strategy, we propose an importance-aware bandwidth-adaptive communication quantization approach, as detailed in Algorithm 1. This method prioritizes the transmission of important parameters with higher precision while quantizing less important parameters have lower precision, ensuring that the total transmitted bits of client $k$ do not exceed the bandwidth budget $M_{k}$.
V-B1 Quantization Levels and Bit Allocation
The available precision levels are predefined as $Q=[32,16,8,4],$ representing the number of bits per parameter in the rank-1 matrix. Specifically, 32-bit represents full precision without quantization, while 16-bit, 8-bit, and 4-bit correspond to integer quantization (INT16, INT8, INT4). The total number of available quantization levels is denoted as $|Q|$. For each quantization level pair $(Q[i],Q[i+1])$, the algorithm computes the bit cost per rank-1 matrix at both higher precision level $b_{\text{high}}$ and the lower precision level $b_{\text{low}}$.
V-B2 Base Cost Calculation
The base cost assumes all rank-1 matrices are transmitted at the lower precision $b_{\text{low}}\cdot N_{k}$, where $N_{k}$ represents the total number of rank-1 matrices for client $k$. For example, $N_{k}=r_{k}$ in the truncation scheme or $N_{k}=(1-\alpha_{k})r_{\text{max}}$ in the freezing scheme. If the bandwidth budget $M_{k}$ is insufficient to cover this base cost, the algorithm skips to the next quantization level pair.
V-B3 High-Precision Allocation
For the available bandwidth $M_{k}$, the algorithm determines the number of rank-1 matrices $n_{h}$ that can be transmitted at the higher precision $Q[i]$, prioritizing matrices with smaller $i$ values (higher importance). The remaining matrices $N_{k}-n_{h}$ are transmitted at the lower precision $Q[i+1]$. Here, the floor function $\lfloor$.$\rfloor$ . ⌋ is used to ensure that $n_{h}$ is an integer.
V-B4 Fallback to Lowest Precision
If no suitable quantization level pair satisfies the bandwidth constraint, the algorithm transmits the most important rank-1 matrices at the lowest available precision $Q[|Q|-1]$. The number of matrices transmitted is determined by the available bandwidth, and any remaining matrices are discarded if the bandwidth is insufficient.
Algorithm 1 Importance-Aware Bandwidth-Adaptive Communication Quantization
0:  LoRA parameters $\mathbf{B}_{k}^{\text{T}},\mathbf{A}_{k}^{\text{T}}$ for each client $k\in\mathcal{K}$ in truncation scheme or $\mathbf{B}_{k}^{\mathcal{I}},\mathbf{A}_{k}^{\mathcal{I}}$ in freezing scheme, total number of rank-1 matrices $N_{k}$, element count per rank-1 matrix $e=d+l$, maximum number of transmit bits $M_{k}$, quantization levels $Q=[32,16,8,4]$.
1:  for $i=0$ to $|Q|-2$ do
2:     $b_{\text{high}}=Q[i]\cdot e$  // Bits per rank-1 matrix at higher precision
3:     $b_{\text{low}}=Q[i+1]\cdot e$  // Bits per rank-1 matrix at lower precision
4:     $b_{\text{base}}=b_{\text{low}}\cdot N_{k}$  // Cost if all at lower precision
5:     if $M_{k}>=b_{\text{base}}$ then
6:        $\delta=b_{\text{high}}-b_{\text{low}}$  // Extra cost at higher precision
7:        $n_{h}=\min\left(N_{k},\left\lfloor\frac{M_{k}-b_{\text{base}}}{\delta}\right\rfloor\right)$
8:        if Truncation Scheme then
9:           Transmit the first $n_{h}$ rank-1 matrices, represented as $\mathbf{B}_{k}^{\text{T}}[:,j]\mathbf{A}_{k}^{\text{T}}[j,:]$ for $j\in[0,n_{h})$, at $Q[i]$-bit precision, and the remaining rank-1 matrices at $Q[i+1]$-bit precision.
10:        else if Freezing Scheme then
11:           Transmit the first $n_{h}$ rank-1 matrices, represented as $\mathbf{B}_{k}^{\mathcal{I}}[:,j]\mathbf{A}_{k}^{\mathcal{I}}[j,:]$ for $j\in[0,n_{h})$, at $Q[i]$-bit precision, and the remaining rank-1 matrices at $Q[i+1]$-bit precision.
12:        end if
13:     end if
14:  end for
15:  $b_{\text{min}}\leftarrow Q[|Q|-1]\cdot e$
16:  $n_{\text{trans}}\leftarrow\left\lfloor\frac{M_{k}}{b_{\text{min}}}\right\rfloor$
17:  Transmit the first $n_{\text{trans}}$ rank-1 matrices at $Q[|Q|-1]$-bit, and discard the remaining rank-1 matrices.
V-C Quantization at Client and Dequantization at Server
In our proposed approach, each LoRA rank-1 matrix is represented as $\mathbf{b}_{i}\mathbf{a}_{i}$, where $\mathbf{b}_{i}\in\mathbb{R}^{d\times 1}$ and $\mathbf{a}_{i}\in\mathbb{R}^{1\times l}$. Depending on the parameter uploading scheme, these rank-1 matrices correspond to either $\mathbf{B}_{k}^{\text{T}}[:,i]\mathbf{A}_{k}^{\text{T}}[i,:]$ in the truncation scheme or $\mathbf{B}_{k}^{\mathcal{I}}[:,i]\mathbf{A}_{k}^{\mathcal{I}}[i,:]$ in the freezing scheme. To reduce communication costs, we quantize vectors $\mathbf{a}_{i}$ and $\mathbf{b}_{i}$ independently using their respective scale and zero point parameters[29]. The quantized vectors and their parameters are then transmitted to the server for dequantization.
V-C1 Quantization at Client
For a given vector $\mathbf{v}$ (which can be either $\mathbf{a}_{i}$ or $\mathbf{b}_{i}$), the quantization process is as follows:
Calculate Scale and Zero Point
For a specified bit quantization precision $q$, the scale $s_{v}$ and zero point $z_{v}$ are computed as follows:
$s_{v}$ $=\frac{\max(\mathbf{v})-\min(\mathbf{v})}{2^{q}-1},$ (33)
$z_{v}$ $=\left\lfloor-\frac{\min(\mathbf{v})}{s_{v}}\right\rceil,$ (34)
where $\max(\mathbf{v})$ and $\min(\mathbf{v})$ represent the maximum and minimum values of the elements in the vector $\mathbf{v}$, respectively.
Quantize the Vector
$\mathbf{v}_{\text{q}}=\text{clamp}\left(\left\lfloor\frac{\mathbf{v}}{s_{v}}+z_{v}\right\rceil,0,2^{q}-1\right),$ (35)
where clamp ($\cdot$) constrains the value within the integer range. Note that the scale $s_{v}$ and zero point $z_{v}$ are scalar values applied element-wise to the vector $\mathbf{v}$, which means each element of $\mathbf{v}$ is quantized independently.
This process is applied separately to $\mathbf{b}_{i}$ and $\mathbf{a}_{i}$, resulting in quantized vectors $\mathbf{b}_{i,\text{q}}$ and $\mathbf{a}_{i,\text{q}}$, along with their respective parameters $s_{b_{i}},z_{b_{i}}$ and $s_{a_{i}},z_{a_{i}}$. The quantized vectors, together with their respective scales and zero points, are then transmitted to the cloud server.
V-C2 Dequantization at Cloud Server
Upon receiving the quantized vectors and respective scales and zero points, the server performs dequantization first to recover approximations of the original vectors:
$\mathbf{b}_{i,\text{dq}}$ $=s_{b_{i}}\cdot(\mathbf{b}_{i,\text{q}}-z_{b_{i}}),$ (36)
$\mathbf{a}_{i,\text{dq}}$ $=s_{a_{i}}\cdot(\mathbf{a}_{i,\text{q}}-z_{a_{i}}),$ (37)
where the dequantization is applied to each quantized rank-1 matrix $\mathbf{b}_{i,\text{q}}$ and $\mathbf{a}_{i,\text{q}}$ using their respective scaling factors $s_{a_{i}},s_{b_{i}}$ and zero points $z_{a_{i}},z_{b_{i}}$.
VI Adaptive Global Aggregation at the Cloud